OpenStudy (anonymous):

The total snowfall in Buffalo, NY one year was 120 inches. If the total snowfall the following year was 90 inches, what was the percent change? Show and explain your solution

OpenStudy (shane_b):

\[100-(\frac{following\space year}{first\space year}*100)=100-(\frac{90}{120}*100)=?\%\space decrease\]

OpenStudy (anonymous):

okay thanks !

OpenStudy (shane_b):

np :)

OpenStudy (anonymous):

:)

OpenStudy (shane_b):

In this one you could also just look at it since the numbers work out perfectly. It went from 120 to 90...so it went from 4/4 to 3/4...a drop of 1/4th which is a 25% decrease.

OpenStudy (anonymous):

Kk tht really helped :)